# Coarse-graining nonequilibrium diffusions with Markov chains

### v1

@@ 1. Introduction
Many physical processes can be described by the interplay of deterministic forcing and random fluctuations. Such processes are typically described as diffusive and have been used to model a plethora of systems including Brownian motion [brown1828brownian], cell motility [Li2011dicty], climate dynamics [Franzke_O’Kane_2017], neural systems [harrison2005stochasticneuronal], and financial markets, amongst others. Mathematically, they can be formulated as Îto stochastic differential equations (SDE), taking the form,
@@
where $X(t)$ represents the state of the process at time $t$, $f(x)$ represents the deterministic forcing, and $\Sigma(x)$ determines the random fluctuations as $W(t)$ is the standard Wiener process [pavliotis2014stochproc].
@@
It has been argued by Schrödinger that biological systems stave off decay by burning energy to maintain themselves in a nonequilibrium steady-state (NESS) [Schrodinger1944whatislife], which has since led to the development of a range of techniques for analysing the properties of a NESS directly from observed, stochastic trajectories of a diffusion [battle2016brokendetailedbalance, Gnesotto2018brokendetailedbalance, Frishman2020stochasticforce, Li2019quantifying]. Central to many of these techniques, is a discretisation of phase-space into discrete voxels, where the evolution of a trajectory in a continuum can be approximated as a sequence of discrete states, as shown in Fig. 1. These methods, and their derivatives, have been used to analyse the nonequilibrium dynamics of many biological systems [battle2016brokendetailedbalance, Gnesotto2018brokendetailedbalance, lynn2021detailedbalance, nartallokaluarachchi2024decomposing]. Nevertheless, such approaches represent a coarse-graining, where much of the underlying continuous information is lost. More specifically, it is known that coarse-graining can obscure the properties of a NESS in a complex, and somewhat unpredictable, manner [Esposito2012coarsegraining]. For example, coarse-graining in both space and time implies that the entropy produced can only be lower bounded from observations of the coarse-grained dynamics [Esposito2012coarsegraining]. Moreover, coarse-graining introduces memory effects, rendering the dynamics non-Markovian [seifert2019inference]. In turn, Markovian estimates of the entropy produced are also lower-bounds on the true value [Schwarz2024Memory]. Within statistical mechanics, coarse-graining has long been a powerful technique to simplify and reduce complex models through techniques like the renormalisation group [Wilson1983renormalisation], or hydrodynamics [Marchetti2013hydrodynamics]. In fact, the stochastic formulation of Brownian motion is itself an abstraction that is obtained by coarse-graining molecular dynamics at a finer scale [Schilling2022coarse].
@@ 2.1. Diffusions in continuous-space: Stochastic differential equations and the Fokker-Planck equation
A continuous-space diffusion, $\{X(t)\in\mathbb{R}^{d}:t>0\}$, is defined as the solution to the stochastic differential equation (SDE),
@@
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a vector field that defines the deterministic forcing, known as the drift, and $D(x)=\frac{1}{2}\Sigma(x)\Sigma^{\top}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is the diffusion matrix, which is positive definite [pavliotis2014stochproc]. The density of the process, $p(x,t)$, is defined as the probability density of the process attaining a value of $x\in\mathbb{R}^{d}$ at time $t>0$. Its dynamics are given by the Fokker-Planck (FP) equation,
@@
Diffusions in a NESS admit the Helmholtz-Hodge decomposition (HHD),
@@
where $f_{\text{rev}}$ and $f_{\text{irr}}$ represent the time-reversible and time-irreversible drift respectively [DaCosta_2023]. This implies that they are even and odd under time-reversal respectively.
@@
In keeping with the second law of thermodynamics, NESS thus have time-irreversible dynamics that lead to the production of entropy. The entropy production rate (EPR) is a key quantity determining the distance of a stationary process from thermodynamic equilibrium. For a diffusion in a NESS, it is given by,
@@
which is only zero if the process is in a ESS, and is positive otherwise [Jiang2004noneq]. Using the HHD, the EPR can be alternatively written as,
@@
which makes it clear that the EPR only depends on the irreversibility of the process, a central result in stochastic thermodynamics [seifert2012thermodynamics].
The condition for stationarity is that $\nabla\cdot J_{\text{ss}}=0$. In a single dimension, this implies that the stationary probability flux must satisfy,
(11) $\frac{d}{dx}(J_{\text{ss}})=0,$
i.e. $J_{\text{ss}}$ is constant over $\mathbb{R}$. In order to guarantee stationarity in unbounded domains, we require an additional asymptotic boundary condition that the probability flux decays at infinity, therefore, for stationary diffusions over $\mathbb{R}$, $J_{\text{ss}}=0$ and the process must be in equilibrium.
Finally, we define a reversible Smoluchowski dynamics to be the particular case where,
(12) $f(x)$ $=-\nabla U(x),$
(13) $D$ $=\frac{\sigma^{2}}{2}I,$
i.e where $f$ is the gradient of a smoothing and containing potential, $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and $D$ is a multiple of the identity. Such diffusion is known as isotropic, and is defined by the noise intensity $\sigma$. In this case, the stationary density is given by,
(14) $\pi(x)$ $=\frac{1}{Z}\exp\left(-\frac{2U(x)}{\sigma^{2}}\right),$
known as a Boltzmann distribution, where $Z$ is a normalising factor, and the noise intensity, $\sigma$ plays the role of temperature [pavliotis2014stochproc]. Although Smoluchowski processes converge to an ESS, any ergodic process with isotropic diffusion has a Boltzmann distribution [^1] To show this, we take $U(x)=-\frac{\sigma^{2}}{2}\log\left(\frac{\pi(x)}{Z}\right)$. Substituting this potential into Eq. (14) returns $\pi(x)$. Moreover, $\nabla U$ is precisely the reversible component of the dynamics from the HHD., including those that converge to a NESS.
@@ 2.2. Discrete-state processes: Continuous-time Markov chains
where $L$ is the Laplacian matrix [Gardiner2010Methods]. The off-diagonal entries, $L_{ij}\geq 0$, represent the transition intensity from states $j$ to $i$. The diagonal entries are then defined to be $L_{ii}=-\sum_{j\neq i}L_{ji}$, which enforces that $\sum_{i}p_{i}(t)=1$ as $\frac{d}{dt}(\sum_{i}p_{i}(t))=0$. Moreover, the waiting distribution in each state is defined to be $\text{Exp}(-L_{ii})$ [Masuda2017randomwalks].
If a CTMC is irreducible and aperiodic, then it is ergodic and will converge to a unique stationary distribution (see Ref. [Ross2019probmodels] for definitions), which satisfies $L\pi=0$. We can define a discrete counterpart to the probability flux vector field,
@@
thus the steady-state is in equilibrium if and only if the flux between all pairs of states vanishes [Jiang2004noneq]. This is more commonly known as the detailed balance condition,
@@
When this condition is violated, the EPR of the NESS can be computed with the Schnakenberg formula [^3] At first glance, the discrete and continuous forms of the EPR may not appear to have much in common. However, both are defined as the relative entropy between the forward and backward path-space measures. This can be derived both measure-theoretically, and more heuristically by taking the derivative of the Shannon entropy of either the distribution, in the case of the CTMC, or of the entropy along a stochastic trajectory, in the case of the diffusion. For further details see Refs. [DaCosta_2023, Jiang2004noneq]. [schnakenberg1976network],
@@
Similarly, the dynamics of the steady-state are reversible in an ESS and irreversible in a NESS [roldan2014thesis].
@@ 3.1. Non-Markovian approaches
Coarse-graining a Markovian diffusion with discrete-states leads to both memory effects and non-Markovian dynamics, in the same way that coarse-graining a Markov chain via state-aggregation induces memory in discrete-state dynamics [Schwarz2024Memory, faccin2021stateaggregations]. This typically leads to both non-exponential waiting times, as well as non-Markovian transitions [^4] It is important to note that these are separate phenomena. Whilst non-exponential waiting times imply a semi-Markov process, coarse-graining can also induce more disruptive memory effects where transition intensities between states depend on the trajectory at large.. Previous approaches have used first-passage time analysis and asymptotic expansions in an attempt to obtain the waiting-time distributions of a discrete-state approximation of prototypical, univariate diffusions [Gernert2014waiting, Falasco2021localdetailed, Ghosala2022inferring, Meyberg2024entropy]. Such an analysis is closely linked to approaches for estimating the EPR from semi-Markov processes [^5] A semi-Markov process is a discrete-state process with arbitrary waiting-time distributions [Ross2019probmodels]. using waiting-time distributions [martinez2019inferring, Skinner2021estimating]. Whilst such approaches more accurately describe the coarse-grained dynamics, they are typically limited in the complexity of the dynamics that can be applied to. Moreover, the statistical inference of semi-Markov processes from experimental data is more involved, hence data analytic methods typically use a Markovian model [battle2016brokendetailedbalance, lynn2021detailedbalance, nartallokaluarachchi2024decomposing]. In discrete-time, it is simple to infer higher-order Markov models from sequence data, and thus reveal the effects of memory in real stochastic dynamics [rosvall2014memory, Sahasrabuddhe2025concise].
Other related works include the coarse-graining of a microscopic Markov process with an SDE in the continuum, or ‘diffusion’, limit, where the convergence of the EPR has also been studied [busiello2019coarsegrained].
@@ 3.2. A Markovian approximation
which leads to a diagonal diffusion matrix with entries,
@@
This, in turn, leads to a FP equation of the form,
@@
where the flux field is given by,
@@
It remains to derive expressions for the numerical flux which ultimately leads to a ME of the form,
@@
A numerical scheme yields a valid ME if $L_{(i,j),(k,l)}>0$ for all off-diagonal entries ($(k,l)\neq(i,j)$), and probability is conserved i.e. the columns of $L$ sum to 0. In Ref. [strang2020applications], Strang considered the centred-difference approximation for the FP equation in 1D, which only yields a valid ME under step-size conditions. Here, we consider the structure-preserving Scharfetter-Gummel discretisation [Frensley2004scharfettergummel], which alleviates these conditions.
@@ 3.3. The Scharfetter-Gummel discretisation
with analogous expressions for the remaining flux terms [Frensley2004scharfettergummel] (for a derivation, see App. A). This naturally leads to a ME equation of the form,
@@
with off-diagonal transition intensities given by,
@@
which are necessarily positive [^6] In cases, where the denominator vanishes, which occurs when the drift vanishes, then we can re-derive the SG discretisation to obtain $L_{(i,j),(i+1,j)}=D^{x}_{i+\frac{1}{2},j}/\Delta x^{2}$, and analogous terms for the other transitions [Frensley2004scharfettergummel].. Moreover, the diagonal intensities are given by,
@@
It is simple to check that the sum,
@@
by replacing $i$ by $i\pm 1$ and $j$ by $j\pm 1$ into the derived intensities. As a result, this scheme yields a valid ME as it has non-negative intensities between states and conserves probability.
It is well-known that FVA conserve probability777i.e. the numerical solution at every time integrates to 1. This is equivalent to the derived Laplacian having columns that sum to 0. [LeVeque2007finitedifference]. Moreover, the SG discretisation preserves the stationary distribution and many thermodynamic properties of the continuous solution [Schlichting2022Scharfetter], hence it is well-suited to this application.
@@ 3.3.1. Entropy production in the Scharfetter-Gummel discretisation
In this section, we will show that the EPR of the discrete approximation converges to that of the underlying diffusion process. We begin by assuming that the discrete process is stationary with distribution $\pi$. As each cell only receives transitions from the four adjacent cells, the EPR of Eq. (18) reduces to,
@@
where each contribution is defined by,
@@
with analogous expressions for the remaining terms. Substituting in the expressions derived from the SG discretisation, we obtain,
@@
Next, we use the approximation $\pi_{i+1,j}=\pi_{i,j}+\Delta x[\partial_{x}\pi]_{i,j}+O(\Delta x^{2})$ and perform a Taylor expansion, which yields,
@@
Using that the stationary distribution will converge i.e. $\pi_{i,j}=\Delta x\Delta y\pi(x_{i},y_{i})+O(\Delta x\Delta y)$, where $\pi(x,y)$ is the stationary density of the diffusion, we have that,
@@
where $J$ is the probability flux associated with the stationary density $\pi$. Thus, the EPR is,

### v2

@@ 1. Introduction
Many physical processes can be described by the interplay of deterministic forcing and random fluctuations. Such processes are typically described as diffusive and have been used to model a plethora of systems including Brownian motion [7], cell motility [41], climate dynamics [Franzke_O’Kane_2017], neural systems [29], and financial markets, amongst others. Mathematically, they can be formulated as Îto stochastic differential equations (SDE), taking the form
@@
where $X(t)$ represents the state of the process at time $t$, $f(x)$ represents the deterministic forcing, and $\Sigma(x)$ determines the random fluctuations as $W(t)$ is the standard Wiener process [58].
@@
It has been argued by Schrödinger that biological systems stave off decay by burning energy to maintain themselves in a nonequilibrium steady-state (NESS) [65], which has since led to the development of a range of techniques for analysing the properties of a NESS directly from observed, stochastic trajectories of a diffusion [2, 26, 20, 40]. Central to many of these techniques, is a discretisation of phase-space into discrete voxels, where the evolution of a trajectory in a continuum can be approximated as a sequence of discrete states, as shown in Fig. 1. These methods, and their derivatives, have been used to analyse the nonequilibrium dynamics of many biological systems [2, 26, 42, 53, 56, 36]. Nevertheless, such approaches represent a coarse-graining, where much of the underlying continuous information is lost. More specifically, it is known that coarse-graining can obscure the properties of a NESS in a complex, and somewhat unpredictable, manner [13]. For example, coarse-graining in both space and time implies that the entropy produced can only be lower bounded from observations of the coarse-grained dynamics [13]. Moreover, coarse-graining introduces memory effects, rendering the dynamics non-Markovian [68]. In turn, Markovian estimates of the entropy produced are also lower-bounds on the true value [66]. Within statistical mechanics, coarse-graining has long been a powerful technique to simplify and reduce complex models through techniques like the renormalisation group [78], or hydrodynamics [43]. In fact, the stochastic formulation of Brownian motion is itself an abstraction that is obtained by coarse-graining molecular dynamics at a finer scale [63].
@@ 2.1. Diffusions in continuous-space: Stochastic differential equations and the Fokker-Planck equation
A continuous-space diffusion, $\{X(t)\in\mathbb{R}^{d}:t>0\}$, is defined as the solution to the stochastic differential equation (SDE)
@@
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a vector field that defines the deterministic forcing, known as the drift, and $D(x)=\frac{1}{2}\Sigma(x)\Sigma^{\top}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is the diffusion matrix, which is positive definite [58]. When using spatially-dependent diffusion, we will use the Îto convention for stochastic integration [58]. [^1] It is worth noting that the choice of convention for stochastic integration can lead to important differences in nonequilibrium properties like time-reversibility [76, 1]. The density of the process, $p(x,t)$, is defined as the probability density of the process attaining a value of $x\in\mathbb{R}^{d}$ at time $t>0$. Its dynamics are given by the Fokker-Planck (FP) equation
@@
Diffusions in a NESS admit the Helmholtz-Hodge decomposition (HHD)
@@
where $f_{\text{rev}}$ and $f_{\text{irr}}$ represent the time-reversible and time-irreversible drift respectively [12]. This implies that they are even and odd under time-reversal respectively.
@@
In keeping with the second law of thermodynamics, NESS thus have time-irreversible dynamics that lead to the production of entropy. The entropy production rate (EPR) is a key quantity determining the distance of a stationary process from thermodynamic equilibrium. For a diffusion in a NESS, it is given by
@@
which is only zero if the process is in a ESS, and is positive otherwise [34]. Using the HHD, the EPR can be alternatively written as
@@
which makes it clear that the EPR only depends on the irreversibility of the process, a central result in stochastic thermodynamics [67].
@@ 2.2. Discrete-state processes: Continuous-time Markov chains
where $L$ is the Laplacian matrix [22]. The off-diagonal entries, $L_{ij}\geq 0$, represent the transition rates from states $j$ to $i$. The diagonal entries are then defined to be $L_{ii}=-\sum_{j\neq i}L_{ji}$, which enforces that $\sum_{i}p_{i}(t)=1$ as $\frac{d}{dt}(\sum_{i}p_{i}(t))=0$. Moreover, the waiting distribution in each state is defined to be $\text{Exp}(-L_{ii})$ [45].
If a CTMC is irreducible and aperiodic, then it is ergodic and will converge to a unique stationary distribution (see Ref. [60] for definitions), which satisfies $L\pi=0$. We can define a discrete counterpart to the probability flux vector field
@@
thus the steady-state is in equilibrium if and only if the flux between all pairs of states vanishes [34]. This is more commonly known as the detailed balance condition
@@
When this condition is violated, the EPR of the NESS can be computed with the Schnakenberg formula [^3] At first glance, the discrete and continuous forms of the EPR may not appear to have much in common. However, both are defined as the relative entropy between the forward and backward path-space measures. This can be derived both measure-theoretically, and more heuristically by taking the derivative of the Shannon entropy of either the distribution, in the case of the CTMC, or of the entropy along a stochastic trajectory, in the case of the diffusion. For further details see Refs. [12, 34]. [32],
@@
Similarly, the dynamics of the steady-state are reversible in an ESS and irreversible in a NESS [59].
@@ 3.1. Non-Markovian approaches
Coarse-graining a Markovian diffusion with discrete-states leads to both memory effects and non-Markovian dynamics, in the same way that coarse-graining a Markov chain via state-aggregation induces memory in discrete-state dynamics [66, 14]. This typically leads to both non-exponential waiting times, as well as non-Markovian transitions [^4] It is important to note that these are separate phenomena. Whilst non-exponential waiting times imply a semi-Markov process, coarse-graining can also induce more disruptive memory effects where transition rates between states depend on the trajectory at large.. Previous approaches have used first-passage time analysis and asymptotic expansions in an attempt to obtain the waiting-time distributions of a discrete-state approximation of prototypical, univariate diffusions [23, 15, 25, 47]. Such an analysis is closely linked to approaches for estimating the EPR from semi-Markov processes [^5] A semi-Markov process is a discrete-state process with arbitrary waiting-time distributions [60]. using waiting-time distributions [44, 71]. Whilst such approaches more accurately describe the coarse-grained dynamics, they are typically limited in the complexity of the dynamics that can be applied to. Moreover, the statistical inference of semi-Markov processes from experimental data is more involved, hence data analytic methods typically use a Markovian model [2, 42, 53]. In discrete-time, it is simple to infer higher-order Markov models from sequence data, and thus reveal the effects of memory in real stochastic dynamics [61, 62].
Other related works include the coarse-graining of a microscopic Markov process with an SDE in the continuum, or ‘diffusion’, limit, where the convergence of the EPR has also been studied [10].
@@ 3.2. A Markovian approximation
which leads to a diagonal diffusion matrix with entries
@@
This, in turn, leads to a FP equation of the form
@@
where the flux field is given by
@@
It remains to derive expressions for the numerical flux which ultimately leads to a ME of the form
@@
A numerical scheme yields a valid ME if $L_{(i,j),(k,l)}>0$ for all off-diagonal entries ($(k,l)\neq(i,j)$), and probability is conserved i.e. the columns of $L$ sum to 0. In Ref. [72], Strang considered the centred-difference approximation for the FP equation in 1D, which only yields a valid ME under step-size conditions. As a result, the problem of deriving a valid approximation is closely related to that of developing a structure-preserving scheme for the FP equation [31, 30]. We will adopt Scharfetter-Gummel discretisation [18], which alleviates step-size conditions.
@@ 3.3. The Scharfetter-Gummel discretisation
with analogous expressions for the remaining flux terms [18] (for a derivation, see App. A). This naturally leads to a ME equation of the form
@@
with off-diagonal transition rates given by
@@
which are necessarily positive [^6] In cases, where the denominator vanishes, which occurs when the drift vanishes, then we can re-derive the SG discretisation to obtain $L_{(i,j),(i+1,j)}=D^{x}_{i+\frac{1}{2},j}/\Delta x^{2}$, and analogous terms for the other transitions [18].. Moreover, the diagonal rates are given by
@@
It is simple to check that the sum
@@
by replacing $i$ by $i\pm 1$ and $j$ by $j\pm 1$ into the derived rates. As a result, this scheme yields a valid ME as it has non-negative rates between states and conserves probability.
It is well-known that FVA conserve probability777i.e. the numerical solution at every time integrates to 1. This is equivalent to the derived Laplacian having columns that sum to 0. [39]. Moreover, the SG discretisation preserves the stationary distribution and many thermodynamic properties of the continuous solution [64], hence it is well-suited to this application.
@@ 3.3.1. Entropy production in the Scharfetter-Gummel discretisation
In this section, we will show that the EPR of the discrete approximation converges to that of the underlying diffusion process. We begin by assuming that the discrete process is stationary with distribution $\pi$. As each cell only receives transitions from the four adjacent cells, the EPR of Eq. (14) reduces to
@@
where each contribution is defined by
@@
with analogous expressions for the remaining terms. Substituting in the expressions derived from the SG discretisation, we obtain
@@
Next, we use the approximation $\pi_{i+1,j}=\pi_{i,j}+\Delta x[\partial_{x}\pi]_{i,j}+O(\Delta x^{2})$ and perform a Taylor expansion, which yields
@@
Using that the stationary distribution will converge i.e. $\pi_{i,j}=\Delta x\Delta y\pi(x_{i},y_{i})+O(\Delta x\Delta y)$, where $\pi(x,y)$ is the stationary density of the diffusion, we have that
@@
where $J$ is the probability flux associated with the stationary density $\pi$. Thus, the EPR is
@@
3.3.2. Recovering the drift and diffusion coefficients
Given the Laplacian matrix, $L$, we can recover the drift and diffusion at discrete points in the grid. To show this explicitly, we consider the rates
(36) $L_{(i,j),(i+1,j)}$ $=\frac{f^{x}_{i+\frac{1}{2},j}e^{-\xi}}{\Delta x(1-e^{-\xi})},$ $L_{(i+1,j),(i,j)}=\frac{f^{x}_{i+\frac{1}{2},j}}{\Delta x(1-e^{-\xi})},$
where $\xi=f^{x}_{i+\frac{1}{2},j}\Delta x/D^{x}_{i+\frac{1}{2},j}$. These are the transition rates from $(i+1,j)$ to $(i,j)$ and its reverse, respectively. We momentarily drop the subscript for convenience, and take the difference in the rates to obtain,
(37) $L_{(i+1,j),(i,j)}-L_{(i,j),(i+1,j)}$ $=\frac{f^{x}-f^{x}e^{-\xi}}{\Delta x(1-e^{-\xi})}=\frac{f^{x}}{\Delta x},$
thus we recover the drift with
(38) $f^{x}_{i+\frac{1}{2},j}$ $=\Delta x\left(L_{(i+1,j)(i,j)}-L_{(i,j)(i+1,j)}\right).$
The remaining drift terms can be obtained analogously. To recover the diffusion, we take the log-ratio of the rates to obtain
(39) $\xi=\log\left(\frac{L_{(i+1,j)(i,j)}}{L_{(i,j)(i+1,j)}}\right)=\frac{f^{x}\Delta x}{D^{x}}.$
We then use our previous expression for the drift to write the diffusion as
(40) $D^{x}_{i+\frac{1}{2},j}$ $=\frac{\Delta x^{2}\left(L_{(i+1,j)(i,j)}-L_{(i,j)(i+1,j)}\right)}{\log\left(L_{(i+1,j)(i,j)}/L_{(i,j)(i+1,j)}\right)}.$
The remaining diffusion coefficients can be recovered in a similar way.
Remark 3.1 (A variational perspective on convergence from Markov chains to diffusions).
It is not immediately obvious that the NESS properties of a CTMC converge to that of the SDE, even when the CTMC is derived from the drift and diffusion of the SDE. In the former, fluctuations have Poissonian statistics stemming from the exponentially-distributed waiting times between transitions, whilst in the latter the fluctuations are driven by Gaussian noise. In order to develop further intuition on this convergence, we consider the variational approach to characterising a NESS, which combines macroscopic fluctuation theory [4], and the large deviation principle [75].
It is useful to define the dissipation potential and its complex conjugate $(\Psi,\Psi^{*})$, which capture the thermodynamic cost of observing a given fluctuation in the flux $J$. Moreover, at steady-state the EPR is given by $\Phi=\Psi+\Psi^{*}$ [48, 49, 50, 35]. It is well-known that the dissipation potential for a diffusive process is quadratic in the flux, leading to the characteristic quadratic form for the EPR, as in Eq. (9) and (10). This implies a ‘linear response’, which, in turn, arises from the Gaussian nature of the fluctuations. On the other hand, the dissipation potential of a CTMC is non-quadratic [48, 35]. However, as discussed and illustrated in App. C, in the continuum limit the potential, and therefore the EPR, is approximately quadratic in the flux. This argument provides a further intuition for the convergence in Sec. 3.3.1.